Circular No. 8889

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET 17P/HOLMES D. Schleicher, Lowell Observatory, obtained eight sets of narrowband

photometry of comet 17P/Holmes on Nov. 1 (r = 2.47 AU) using the Hall 1.1-m telescope at Lowell Observatory. Projected aperture radii ranged from 14500 to 120200 km, and derived production rates exhibited large trends with aperture size for all species — with the highest values always at the largest apertures, consistent with a slowly decreasing rate of overall activity since the initial outburst. Production-rate ranges for each species are: log Q(OH) = 29.60–29.80; equivalent log Q(water; vectorial) = 29.54– 29.74; log Q(NH) = 27.65–27.89; log Q(CN) = 27.26–27.51; log Q(C_{2}) = 27.33–27.61; log Q(C_{3}) = 26.64–26.87; log [Afρ] (526 nm) = 5.13–5.74 (cf. IAUC 7342). The resulting abundance ratios indicate that comet 17P has ‘typical’ composition (based on A’Hearn et al. 1995, Icarus 118, 223), but an exceptionally high dust-to-gas ratio, possibly associated with the finite lifetimes of the gas molecules released at the onset of the outburst as compared to long-lived dust grains. A significant contribution to relatively high dust abundance may also be caused by a portion of the dust tail remaining within the photometer apertures due to projection effects from the comet’s small phase angle.

Further naked-eye total-magnitude estimates (cf. IAUC 8887): Oct. 27.63 UT, 2.4 (Y. Nagai, Gunma, Japan); 28.10, 2.5 (J. Rao, Putnam Valley, NY, U.S.A.); 28.75, 2.5 (H. Dahle, Les Olives, France); 29.54, 2.3 (S. Yoshida, Ibaraki, Japan); 30.76, 2.4 (T. Karhula, Vasteras, Sweden); 31.76, 2.5 (K. Cernis, Vilnius, Lithuania); Nov. 1.79, 2.3 (J. J. Gonzalez, Leon, Spain); 2.96, 2.5 (A. Pereira, Cabo da Roca, Portugal).

COMETS C/2007 L12, C/2007 L13, AND C/2007 M4 (SOHO) Further to IAUC 8888, additional Kreutz sungrazing comets have been

found on SOHO website images (“discovery” observations tabulated below). C/2007 L12 was extremely faint (mag ∼ 8.5–9.0). C/2007 L13 and C/2007 M4 were stellar (mag ∼ 5.5) in C3 images. C/2007 L13 appeared slightly diffuse with a very faint tail in C2 images. C/2007 M4 was slightly diffuse and showed a short, thin, faint tail in C2 images.

Comet 2007 UT α_{2000} δ_{2000} Inst. F MPEC
C/2007 L12 June 14.854 5^{h}25.^{m}8 +21^{o}22′ C2  RM 2007-T122
C/2007 L13  15.138 5 26.2 +19 58 C3/2 AK 2007-T122
C/2007 M4  24.696 6 01.0 +19 16 C3/2 TH 2007-U15
2007 November 2  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2007 CBAT  Daniel W. E. Green